# MATRIX AS PLAN: Structured Logical Reasoning with Feedback-Driven Replanning

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologiesccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Natural language processing
@@ 1. Introduction
The Web’s explosive growth has created a vast, heterogeneous substrate of knowledge. Extracting task-relevant content and reliable semantic knowledge from these repositories is essential (Penedo et al., 2024). The necessity has motivated the integration of Large Language Models (LLMs) into web-based applications, including browsing-based question answering, cross-site web agents, and LLM-assisted information extraction and knowledge-graph construction (Nakano et al., 2021; Yoran et al., 2024; Xu et al., 2024a; Chen and Bertozzi, 2023). Despite strong capacity in natural language generation, text classification, and machine translation (Achiam et al., 2023; Liu et al., 2024; Comanici et al., 2025), current LLMs remain unreliable on complex logical reasoning (Shojaee et al., 2025; Roh et al., 2025). The gap is most evident in higher-order tasks such as mathematical problem solving and intricate logical inference. Consequently, to ensure trustworthy behavior in web applications, developing robust reasoning is essential for advancing LLMs within the Web’s semantics and knowledge ecosystem.
@@
To address the above limitations, researchers have explored several neuro-symbolic approaches. One prominent line of work emphasizes a hybrid architecture that integrates an LLM with a symbolic reasoning backend (Gao et al., 2023; Ryu et al., 2024): 1) LLM firstly translates a natural language statement of the reasoning problem into a set of first-order logic formulas; and 2) a symbolic solver uses this formal representation to automatically plan the reasoning steps, execute the derivations, and produce the final answer. Within this pipeline, the symbolic solver plays a pivotal role in both planning and execution. It imposes constraints and performs verification to mitigate reasoning errors that arise from the inherent abstraction and vagueness of natural language formulations (Ye et al., 2023; Pan et al., 2023; Kirtania et al., 2024; Olausson et al., 2023).
However, external reasoners require extremely strict formatting on their inputs. Even slight deviations introduced by an LLM during generation can cause parsing to fail and the entire reasoning pipeline to halt. To fill the gap, LLM-driven reasoning methods (Xu et al., 2024c, b; Liu et al., 2025; Li et al., 2024) are designed to syntactic noise and structural perturbations. A representative example is SymbCoT (Xu et al., 2024c): rather than relying on any symbolic solver, LLM is adopted in a systematic four-stage pipeline: translation, planning, solving, and verification, suggesting that LLM–driven reasoning frameworks offer strong generative flexibility. However, in practice, they exhibit limitations that may degrade reasoning performance:
@@
    Reasoning by LLMs on prose-style text intermixes predicates, entities, and premises within free-form narration. It maintains only loose alignment through inline annotations (e.g., :::). In the absence of a unified namespace and field-level constraints, types and variable scope could not be statically checked, and cross-sentential references are prone to drift. As shown in Fig. 1(a), in an unstructured rendering without a unified namespace, an expression such as $\mathrm{Sell}(y,\mathrm{oneMillionCopies})$ inserts oneMillionCopies as a bare constant in the second argument. This token is neither defined in the predicate inventory nor annotated with type or unit. Consequently, the rule $\forall y\,\big(\mathrm{Sell}(y,\mathrm{oneMillionCopies})\rightarrow\mathrm{Selected}(y)\big)$ can only be treated passively as a “special-case match,” rather than expressing the computable semantics $\text{sales}\geq 10^{6}$. Signature constraints and static checks therefore cannot be enforced. The system cannot promptly or reliably detect and constrain such semantic and typing omissions, nor can it proactively require completion. Subsequent reasoning then runs on incomplete dependencies, making unstable or erroneous outcomes likely. By contrast, structured symbolic translation, under a unified namespace, uses keys and indices for precise addressing. It can perform static checks and consistency validation for type consistency, variable scope, and cross-sentential references. With equivalent informational content, it thus yields higher decidability and overall accuracy.
@@ 2.1. Structured Data Management
Recent studies show that converting the reasoning process into structured representations can alleviate the issues above (Wu et al., 2025). Structured carriers reduce semantic ambiguity and explicitly maintain dependencies between pieces of information. In CoT setting, embedding intermediate reasoning steps into parseable fields significantly improves the accuracy of mathematical and symbolic reasoning. Some reports even claim that adding just a single field to carry the reasoning yields about a 60% gain on GSM8K (Leo, 2024). Going further, the schema design itself directly affects performance (Tam et al., 2024). Small changes in field names or ordering can cause orders-of-magnitude differences; for example, renaming ‘final_choice’ to ‘answer’ can raise accuracy from 4.5% to 95% (Leo, 2024).
Inspired by the above research, we propose a prompt-driven structured representation that covers the full pipeline of symbolic reasoning. For illustration, in symbol translation process, we define a parsing map from a natural language sequence $x_{1:T}$ (a sequence of text spans of length $T$) to symbolic objects:
@@
(3) $\mathrm{enc}(S)=\{(NL_{i},Symbol_{i})|i=1,\dots,N\}.$
This general mapping captures the correspondence between each natural language statement and its canonical symbolic form, facilitating downstream auditing and downstream reasoning. Compared with free-text representations that maintain only a loosely aligned structure via inline markers (e.g., “:::”), this representation reduces semantic ambiguity and the propagation of redundancy. It also provides stable anchors for planners and solvers. In feedback-driven settings, it further supports auditable intermediate evidence and replanning. Without sacrificing expressivity, this structured representation improves the reliability and controllability of complex symbolic reasoning.
@@ 2.3. Feedback-Driven Replanning Mechanism
(9) $f_{\text{diag}}:(\tau,D)\longrightarrow\mathcal{E}.$
Where $\mathcal{E}$ is a set of failure labels. It covers categories such as ‘missing prerequisites,” ‘rule misuse,” ‘premature termination,” and ‘redundancy.” In this way, diagnosis is converted from plain text to a structured object $\bigl(V,A,\mathcal{E}\bigr)$, thus providing auditable anchors for subsequent plan editing and verification.
@@
(11) $\Pi^{\prime\prime}=\mathrm{Normalize}(\Pi^{\prime}).$
Where $\mathrm{Normalize}$ removes redundancy, breaks cycles, and ensures a valid topological ordering. Re-execution depends only on ‘the prior context ++ the currently revised plan.” The process is simple and reproducible:
(12) $\mathrm{Exec}(C,\Pi^{\prime\prime})\longrightarrow(y,T^{\prime}).$
Where $y$ is the new answer and $T^{\prime}$ is the new log. $\Pi^{\prime\prime}$ fills in the necessary prerequisites and removes spurious or transitively redundant dependencies. Normalization ensures that the main chain executes completely in topological order. With these guarantees, re-execution avoids biases such as ‘premature stop,” ‘missing reasoning,” and ‘dependency errors.” Thus, without changing the input context $C$, it yields more accurate and robust answers for complex logical reasoning.
@@ 3.2. Step 2: Matrix-based Planner
From the structured representation $\mathcal{H}=\{P_{c},Q_{c}\}$ given by the symbol translator, a matrix-based planner constructs an executable dependency plan $\Pi=(V,A)$. Here, $V=\{v_{1},\dots,v_{m}\}$ is a finite set of reasoning steps. $A_{ij}\in\{0,1\}^{N\times N}$ is a dependency matrix. In this dependency matrix $A_{ij}$, we stipulate the following rule: if $A_{ij}=1\ \Rightarrow\ v_{i}\prec v_{j}$, then $v_{i}$ must be executed before $v_{j}$. The graph formed by $A$ is required to be a directed acyclic graph, which ensures the existence of a topological order to obtain a valid execution schedule and avoids execution difficulties caused by interleaved dependencies. Accordingly, we obtain a dependency plan $\Pi$ that is aligned with $P_{c}$ and $Q_{c}$ and has a clear structure, thereby balancing dependency and executability without sacrificing expressive power.
@@ 4.1. Experimental Setup
Dataset. We evaluate on the same set of five logical–reasoning datasets: AR-LSAT (Zhong et al., 2022), LogicalDeduction (Srivastava et al., 2023), FOLIO (Han et al., 2022), Proof-
Writer (Tafjord et al., 2020), and PrOntoQA (Saparov and He, 2022). These datasets span multiple symbolic representations and problem configurations, allowing us to test applicability and robustness across heterogeneous settings. PrOntoQA, ProofWriter, and FOLIO follow a first-order logic paradigm. LogicalDeduction and AR-LSAT are predominantly constraint-based. Together they cover different reasoning depths and choice cardinalities (e.g., T/F/U vs. multiple options). To avoid bias from differences in data splits, we adopt the same test subsets as Logic-LM and SymbCoT. During evaluation, any output with formatting noncompliance or missing fields is counted as incorrect. The detailed prompt temples are presented in Appendix A.1.
@@
Baselines. We adopt representative methods as baselines: Logic-LM (Pan et al., 2023), SymbCoT(Xu et al., 2024c), and Aristotle (Xu et al., 2024b). We also include two generic prompting styles—standard prompting and chain-of-thought (CoT) prompting that leverages the in-context learning capability of base LLMs. We conduct a unified evaluation on the five datasets described above. However, to ensure strictly fair comparability, the public implementation of Aristotle provides neither results nor reproducible configurations for FOLIO, AR-LSAT, or LogicalDeduction. Accordingly, we report their numbers only on PrOntoQA and ProofWriter; the remaining three entries are marked with “—”.
@@ 4.1.1. Overall Performance
As shown in Tables LABEL:4o and  LABEL:cross, MatrixCoT attains the best—or tied best—accuracy on both closed-source and open-source model families. Its variance across models and datasets is small. This indicates strong robustness and transferability. To ensure comparability, all methods use identical prompt templates and output constraints on each model. Under this setting, our method still leads. The gains therefore stem primarily from the method itself rather than from tailoring to a particular model or prompt.
@@ 4.3. Analysis and Discussion
Error Analysis. Reliable symbolic reasoning depends on three capabilities: 1) Information acquisition and structuring with target-relevant content; 2) Path selection and plan orchestration by explicitly modeling precedence–successor constraints and maintaining global consistency; and 3) Executing tasks with constraints and accuracy by performing semantically correct derivations according to the specified plan. MatrixCoT strengthens the first two via structured reasoning data management and a matrix-based plan. Structured reasoning data management provides a typed, alignable symbolic base. The matrix plan fixes executable precedence relations through a dependency matrix $A$. Together, they make “finding the right information and laying out the correct path” a high-probability event. On this basis, execution is typically smoother and yields robust gains. For models with weak instruction, execution can still fail. It occurs even when the plan is logically complete and its dependencies are explicit. The failures often stem from an inadequate grasp of the dependency constraints. Thus, MatrixCoT reduces reliance on free-text planning and ad hoc search, but it does not substitute for the base model’s core execution ability. As a model’s basic comprehension and compliance improve with scale, it exploits the matrix plan more effectively, and the gains become more stable.
@@
4.4. Conclusion
In this paper, we introduce MatrixCoT, a structured CoT framework that integrates a matrix-based method and a feedback-driven replanning mechanism. Specifically, we enforce a unified structured output format and normalize solution steps as typed symbolic representations, while explicitly encoding cross-step constraints via a dependency matrix. We also derive a verifiable, executable stepwise plan, and introduce a feedback-driven replanning mechanism, which can repair broken chains and corrects deviations. The entire procedure operates without any external solver, yet improves robustness, interpretability, and controllability. Extensive experiments are conducted on five standard benchmarks and five LLms (both closed and open source), demonstrating stable and competitive performance across all settings.
@@ Acknowledgements.
The above work was supported in part by the Joint Funds of the National Natural Science Foundation of China under Grant U25A20436, Supported by Guangdong S&T Programme (No. 2025B0101120006), the National Natural Science Foundation of China (NSFC) (62372047, 62302049), Guangxi Key Research & Development Program (FN2504240036, 2025FN96441087), the Natural Science Foundation of Guangdong Province (2024A1515011323), the Supplemental Funds for Major Scientific Research Projects of Beijing Normal University, Zhuhai (ZHPT2023002), the Fundamental Research Funds for the Central Universities, Ministry of Education Industry-University Cooperation Collaborative Education Project under 240904497110437, Higher Education Research Topics of Guangdong Association of Higher Education in the 14th Five-Year Plan under 24GYB207. Thank support from the Interdisciplinary Intelligence Super Computer Center of Beijing Normal University at Zhuhai.

### v2

@@ Abstract.
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesisbn: 979-8-4007-2307-0/2026/04doi: 10.1145/XXXXXX.XXXXXXccs: Computing methodologiesccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Natural language processing
@@ 1. Introduction
The Web’s explosive growth has created a vast, heterogeneous substrate of knowledge. Extracting task-relevant content and reliable semantic knowledge from these repositories is essential (Penedo et al., 2024; Liu et al., 2025a; Pan et al., 2025). This necessity has motivated the integration of Large Language Models (LLMs) into web-based applications, including browsing-based question answering, cross-site web agents, and LLM-assisted information extraction and knowledge-graph construction (Nakano et al., 2021; Yoran et al., 2024; Xu et al., 2024a; Chen and Bertozzi, 2023). Despite their strong capabilities in natural language generation, text classification, and machine translation (Achiam et al., 2023; Liu et al., 2024; Comanici et al., 2025), current LLMs remain unreliable on complex logical reasoning (Shojaee et al., 2025; Roh et al., 2025). The gap is most evident in higher-order tasks such as mathematical problem solving and intricate logical inference. Consequently, to ensure trustworthy behavior in web applications, developing robust reasoning is essential for advancing LLMs within the Web’s semantics and knowledge ecosystem.
@@
To address the above limitations, researchers have explored several neuro-symbolic approaches. One prominent line of work emphasizes a hybrid architecture that integrates an LLM with a symbolic reasoning backend (Gao et al., 2023; Ryu et al., 2025): 1) The LLM first translates a natural language statement of the reasoning problem into a set of first-order logic formulas; and 2) a symbolic solver uses this formal representation to automatically plan the reasoning steps, execute the derivations, and produce the final answer. Within this pipeline, the symbolic solver plays a pivotal role in both planning and execution. It imposes constraints and performs verification to mitigate reasoning errors that arise from the inherent abstraction and vagueness of natural language formulations (Ye et al., 2023; Pan et al., 2023; Kirtania et al., 2024; Olausson et al., 2023).
However, external reasoners require extremely strict formatting on their inputs. Even slight deviations introduced by an LLM during generation can cause parsing to fail and the entire reasoning pipeline to halt. To fill the gap, LLM-driven reasoning methods (Xu et al., 2024b, 2025; Liu et al., 2025b) are designed to be robust to syntactic noise and structural perturbations. A representative example is SymbCoT (Xu et al., 2024b): rather than relying on any symbolic solver, the LLM is adopted in a systematic four-stage pipeline: translation, planning, solving, and verification, suggesting that LLM–driven reasoning frameworks offer strong generative flexibility. However, in practice, they exhibit limitations that may degrade reasoning performance:
@@
    Reasoning by LLMs on prose-style text intermixes predicates, entities, and premises within free-form narration. It maintains only loose alignment through inline annotations (e.g., :::). In the absence of a unified namespace and field-level constraints, types and variable scope can not be statically checked, and cross-sentential references are prone to drift. As shown in Fig. 1(a), in an unstructured rendering without a unified namespace, an expression such as $\mathrm{Sell}(y,\mathrm{oneMillionCopies})$ inserts oneMillionCopies as a bare constant in the second argument. This token is neither defined in the predicate inventory nor annotated with a type or a unit. Consequently, the rule $\forall y\,\big(\mathrm{Sell}(y,\mathrm{oneMillion}$$\mathrm{Copies})\rightarrow\mathrm{Selected}(y)\big)$ can only be treated passively as a “special-case match,” rather than expressing the computable semantics $\text{sales}\geq 10^{6}$. Signature constraints and static checks therefore cannot be enforced. The system cannot promptly or reliably detect and constrain such semantic and typing omissions, nor can it proactively require completion. Subsequent reasoning then runs on incomplete dependencies, making unstable or erroneous outcomes likely. By contrast, structured symbolic translation, under a unified namespace, uses keys and indices for precise addressing. It can perform static checks and consistency validation for type consistency, variable scope, and cross-sentential references. With equivalent informational content, it thus yields higher decidability and overall accuracy.
@@ 2.1. Structured Data Management
Recent studies show that converting the reasoning process into structured representations can alleviate the issues above (Wu et al., 2025). Structured carriers reduce semantic ambiguity and explicitly maintain dependencies between pieces of information. In CoT settings, embedding intermediate reasoning steps into parseable fields significantly improves the accuracy of mathematical and symbolic reasoning. Some reports even claim that adding just a single field to carry the reasoning yields about a 60% gain on GSM8K (Leo, 2024). Going further, the schema design itself directly affects performance (Tam et al., 2024). Small changes in field names or ordering can cause orders-of-magnitude differences; for example, renaming “final_choice” to “answer” can raise accuracy from 4.5% to 95% (Leo, 2024).
Inspired by the above research, we propose a prompt-driven structured representation that covers the full pipeline of symbolic reasoning. For illustration, in the symbol translation process, we define a parsing map from a natural language sequence $x_{1:T}$ (a sequence of text spans of length $T$) to symbolic objects:
@@
(3) $\mathrm{enc}(S)=\{(NL_{i},Symbol_{i})\mid i=1,\dots,N\}.$
This general mapping captures the correspondence between each natural language statement and its canonical symbolic form, facilitating downstream auditing and reasoning. Compared with free-text representations that maintain only a loosely aligned structure via inline markers (e.g., “:::”), this representation reduces semantic ambiguity and the propagation of redundancy. It also provides stable anchors for planners and solvers. In feedback-driven settings, it further supports auditable intermediate evidence and replanning. Without sacrificing expressivity, this structured representation improves the reliability and controllability of complex symbolic reasoning.
@@ 2.3. Feedback-Driven Replanning Mechanism
(9) $f_{\text{diag}}:(\tau,D)\longrightarrow\mathcal{E},$
where $\mathcal{E}$ is a set of failure labels. It covers categories such as “missing prerequisites,” “rule misuse,” “premature termination,” and “redundancy.” In this way, diagnosis is converted from plain text to a structured object $\bigl(V,A,\mathcal{E}\bigr)$, thus providing auditable anchors for subsequent plan editing and verification.
@@
(11) $\Pi^{\prime\prime}=\mathrm{Normalize}(\Pi^{\prime}),$
where $\mathrm{Normalize}$ removes redundancy, breaks cycles, and ensures a valid topological ordering. Re-execution depends only on “the prior context ++ the currently revised plan.” The process is simple and reproducible:
(12) $\mathrm{Exec}(C,\Pi^{\prime\prime})\longrightarrow(y,T^{\prime}),$
where $y$ is the new answer and $T^{\prime}$ is the new log. $\Pi^{\prime\prime}$ fills in the necessary prerequisites and removes spurious or transitively redundant dependencies. Normalization ensures that the main chain executes completely in topological order. With these guarantees, re-execution avoids biases such as “premature stop,” “missing reasoning,” and “dependency errors.” Thus, without changing the input context $C$, it yields more accurate and robust answers for complex logical reasoning.
@@ 3.2. Step 2: Matrix-based Planner
From the structured representation $\mathcal{H}=\{P_{c},Q_{c}\}$ given by the symbol translator, a matrix-based planner constructs an executable dependency plan $\Pi=(V,A)$. Here, $V=\{v_{1},\dots,v_{m}\}$ is a finite set of reasoning steps. $A=[A_{ij}]_{i,j=1}^{m}\in\{0,1\}^{m\times m}$ is a dependency matrix. For each entry $A_{ij}$, if $A_{ij}=1$, then $v_{i}\prec v_{j}$, i.e., $v_{i}$ must be executed before $v_{j}$. The graph formed by $A$ is required to be a directed acyclic graph, which ensures the existence of a topological order to obtain a valid execution schedule and avoids execution difficulties caused by interleaved dependencies. Accordingly, we obtain a dependency plan $\Pi$ that is aligned with $P_{c}$ and $Q_{c}$ and has a clear structure, thereby balancing dependency and executability without sacrificing expressive power.
@@ 4.1. Experimental Setup
Dataset. We evaluate on the same set of five logical–reasoning datasets: AR-LSAT (Zhong et al., 2022), LogicalDeduction (Srivastava et al., 2023), FOLIO (Han et al., 2024), Proof-
Writer (Tafjord et al., 2021), and PrOntoQA (Saparov and He, 2023). These datasets span multiple symbolic representations and problem configurations, allowing us to test applicability and robustness across heterogeneous settings. PrOntoQA, ProofWriter, and FOLIO follow a first-order logic paradigm. LogicalDeduction and AR-LSAT are predominantly constraint-based. Together they cover different reasoning depths and choice cardinalities (e.g., T/F/U vs. multiple options). To avoid bias from differences in data splits, we adopt the same test subsets as Logic-LM and SymbCoT. During evaluation, any output with formatting noncompliance or missing fields is counted as incorrect. The detailed prompt templates are presented in Appendix A.1.
@@
Baselines. We adopt representative methods as baselines: Logic-LM (Pan et al., 2023), SymbCoT(Xu et al., 2024b), and Aristotle (Xu et al., 2025). We also include two generic prompting styles—standard prompting and chain-of-thought (CoT) prompting that leverages the in-context learning capability of base LLMs. We conduct a unified evaluation on the five datasets described above. However, to ensure strictly fair comparability, the public implementation of Aristotle provides neither results nor reproducible configurations for FOLIO, AR-LSAT, or LogicalDeduction. Accordingly, we report their numbers only on PrOntoQA and ProofWriter; the remaining three entries are marked with “—”.
4.2. Main Result
@@ 4.2.1. Overall Performance
As shown in Tables 1 and  2, MatrixCoT attains the best—or tied best—accuracy on both closed-source and open-source model families. Its variance across models and datasets is small. This indicates strong robustness and transferability. To ensure comparability, all methods use identical prompt templates and output constraints on each model. Under this setting, our method still leads. The gains therefore stem primarily from the method itself rather than from tailoring to a particular model or prompt.
@@ 4.4. Analysis and Discussion
Error Analysis. Reliable symbolic reasoning depends on three capabilities: 1) Information acquisition and structuring with target-relevant content; 2) Path selection and plan orchestration by explicitly modeling precedence–successor constraints and maintaining global consistency; and 3) Executing tasks with constraints and accuracy by performing semantically correct derivations according to the specified plan. MatrixCoT strengthens the first two via structured reasoning data management and a matrix-based plan. Structured reasoning data management provides a typed, alignable symbolic base. The matrix plan fixes executable precedence relations through a dependency matrix $A$. Together, they make “finding the right information and laying out the correct path” a high-probability event. On this basis, execution is typically smoother and yields robust gains. For models with weak instruction-following ability, execution can still fail. It occurs even when the plan is logically complete and its dependencies are explicit. The failures often stem from an inadequate grasp of the dependency constraints. Thus, MatrixCoT reduces reliance on free-text planning and ad hoc search, but it does not substitute for the base model’s core execution ability. As a model’s basic comprehension and compliance improve with scale, it exploits the matrix plan more effectively, and the gains become more stable.
@@
5. Conclusion
In this paper, we introduce MatrixCoT, a structured CoT framework that integrates a matrix-based method and a feedback-driven replanning mechanism. Specifically, we enforce a unified structured output format and normalize solution steps as typed symbolic representations, while explicitly encoding cross-step constraints via a dependency matrix. We also derive a verifiable, executable stepwise plan, and introduce a feedback-driven replanning mechanism, which can repair broken chains and correct deviations. The entire procedure operates without any external solver, yet improves robustness, interpretability, and controllability. Extensive experiments are conducted on five standard benchmarks and five LLMs (both closed- and open-source), demonstrating stable and competitive performance across all settings.
@@ Acknowledgements.
This work was supported in part by the Joint Funds of the National Natural Science Foundation of China under Grant U25A20436; by the Guangxi Key Research & Development Program (FN2504240036, 2025FN96441087); by the National Natural Science Foundation of China (NSFC) (62372047, 62302049); by the Natural Science Foundation of Guangdong Province (2024A1515011323); by the Supplemental Funds for Major Scientific Research Projects of Beijing Normal University, Zhuhai (ZHPT2023002); by the Fundamental Research Funds for the Central Universities; by the Industry–University Cooperation Collaborative Education Project of the Ministry of Education (240904497110437); and by the Higher Education Research Topics of the Guangdong Association of Higher Education in the 14th Five-Year Plan (24GYB207). We acknowledge the support of the Interdisciplinary Intelligence Super Computer Center of Beijing Normal University at Zhuhai.